Potassium hydrogen phthalate is a solid, monoprotic acid frequently used in the laboratory as a primary standard. It has the unwieldy formula of KHC8H4O4. This is often written in shorthand notation as KHP.

If 34.55 mL of a barium hydroxide solution are needed to neutralize 2.211 grams of KHP, what is the concentration (mol/L) of the barium hydroxide solution?

_______ M
1
Expert's answer
2019-03-12T08:44:00-0400

Ba(OH)2+ 2KHC8H4O4= Ba(KC8H4O4)2+ 2H2O

M(KHC8H4O4) = 39 + 12*8 + 5 + 16*4 = 204.22 (g/mol)

n(KHC8H4O4) = m/M = 2.211 / 204.22 = 0.01 (mol)

n(Ba(OH)2) = 2 * n(KHC8H4O4) = 0.02 mol

c(Ba(OH)2) = n/V = 0.02 / 0.03455 = 0.58 M

Answer: c(Ba(OH)2) = 0.58 M